I am completely lost when it comes to polynomials.
You dive directly upward from a board that is 48 feet high. After "t" seconds, your height above the water is described by the polynomial:
-16t^2 + 32t + 48
a. Factor the polynomial completely. Begin by factoring -16 from each term.
b. Evaluate both the original polynomial and its factored form for t = 3. Do you get the same answer for each evaulation? Describe what this answer means.

Answer by checkley71(8403)

-16t~2+32t+48=0 dividing each term by -16 gives t~2-2t+3=0 or (t-3)(t+1)=0
or t=3 & t=-1. This says you are 3 feet above the board.